how do i work the following problem: (3x-9)(x+1)=0
Answer by checkley71(8403)

SETTING EACH FACTOR TO O GIVES YOU THE TWO SOLUTIONS THUS
3X-9=0 OR 3X=9 OR X=9/3 or x=3 and x+1=0 or x=-1